Payroll Entries

Widmer Company had gross wages of $270,000 during the week ended June 17. The amount of wages subject to social security tax was $243,000, while the amount of wages subject to federal and state unemployment taxes was $34,000. Tax rates are as follows:

Social security 6.0%
Medicare 1.5%
State unemployment 5.3%
Federal unemployment 0.8%

The total amount withheld from employee wages for federal taxes was $54,000.

If an amount box does not require an entry, leave it blank.

a. Journalize the entry to record the payroll for the week of June 17.

Wages Expense
Social Security Tax Payable
Medicare Tax Payable
Employees Federal Income Tax Payable
Wages Payable

Feedback

Gross pay is the amount that employees have earned before taxes and deductions. A portion of employees' earnings are owed for such items as state and federal taxes. Net pay is also known as take home pay.

Learning Objective 2.

b. Journalize the entry to record the payroll tax expense incurred for the week of June 17.

Payroll Tax Expense
Social Security Tax Payable
Medicare Tax Payable
State Unemployment Tax Payable
Federal Unemployment Tax Payable
Answer #1
a
Wages Expense 270000
        Social Security Tax Payable 14580 =243000*6%
        Medicare Tax Payable 4050 =270000*1.5%
        Employees Federal Income Tax Payable 54000
        Wages Payable 197370
b
Payroll Tax Expense 20704
        Social Security Tax Payable 14580 =243000*6%
        Medicare Tax Payable 4050 =270000*1.5%
        State Unemployment Tax Payable 1802 =34000*5.3%
        Federal Unemployment Tax Payable 272 =34000*0.8%